# Analytic and Variational Stability of Deep Learning Systems

### v1

@@ 7.1 Locally Lipschitz learning flows and Clarke derivatives
is locally Lipschitz on $\mathcal{X}_{k}\times\Theta\times\mathcal{U}$. There exist constants
@@
the blocks $V_{x}$, $V_{\theta}$, and $V_{u}$ satisfy
@@
For a locally Lipschitz map $F\colon\mathbb{R}^{m}\to\mathbb{R}^{n}$, the Clarke generalized Jacobian at $z$ is the compact convex set
$\partial_{C}F(z)=\mathrm{co}\Bigl\{\lim_{i\to\infty}JF(z_{i})\;\Big|\;z_{i}\to z,\,F\text{ differentiable at }z_{i}\Bigr\},$
where $JF(z_{i})$ denotes the classical Jacobian and $\mathrm{co}$ denotes convex hull [Clarke1990Optimization, Chapter 2]. Local Lipschitz continuity guarantees that $\partial_{C}F(z)$ is nonempty and upper semicontinuous in $z$ [Clarke1990Optimization, Clarke1998Nonsmooth].
Assumption 4 covers a wide class of non-smooth components used in learning systems. Piecewise affine activations such as ReLU and leaky ReLU are globally Lipschitz. Proximal operators of proper lower semicontinuous convex functionals are firmly nonexpansive and therefore Lipschitz with constant one [RockafellarWets1998]. Orthogonal and convex projections satisfy the same property [BauschkeCombettes2011]. Compositions of these maps with affine layers preserve local Lipschitz continuity and yield finite constants $L^{\mathrm{C}}_{\mathrm{fwd}}$, $L^{\mathrm{C}}_{\mathrm{par}}$, and $L^{\mathrm{C}}_{\mathrm{upd}}$.
Under Assumption 4, the learning dynamics are modelled as a discrete-time difference inclusion
@@
where the multivalued notation reflects possible non differentiability. Existence of trajectories follows from standard results for discrete-time systems with locally Lipschitz right-hand sides [Clarke1990Optimization, Chapter 2].
7.2 Generalized Learning Stability Profile and exponents
We now extend the Learning Stability Profile to the non-smooth setting. Sensitivities are measured using the largest block norms over all elements of the Clarke generalized Jacobian.
@@ 7.2 Generalized Learning Stability Profile and exponents
Under Assumption 4, the generalized stability signature at $(k,t)$ and state $S^{t}_{k}$ is the triple
@@
When $\Phi_{k,t}$ is differentiable at $S^{t}_{k}$, the Clarke generalized Jacobian reduces to the singleton containing the classical Jacobian, and the generalized signature coincides with the smooth Learning Stability Profile. In general, the generalized profile captures the worst-case local amplification of infinitesimal perturbations induced by any active linearisation of the non-smooth learning flow.
@@
$\tilde{\alpha}_{x}$ $:=\limsup_{L,T\to\infty}\frac{1}{L+T}\log\Bigl(\sup_{0\leq t\leq T}\sigma^{x,C}_{k,t}\Bigr),$
$\tilde{\alpha}_{\theta}$ $:=\limsup_{L,T\to\infty}\frac{1}{L+T}\log\Bigl(\sup_{0\leq t\leq T}\sigma^{\theta,C}_{k,t}\Bigr),$
$\tilde{\alpha}_{u}$ $:=\limsup_{L,T\to\infty}\frac{1}{L+T}\log\Bigl(\sup_{0\leq t\leq T}\sigma^{u,C}_{k,t}\Bigr).$
@@
The learning system is generalized analytically stable of order $(\tilde{\beta}_{x},\tilde{\beta}_{\theta},\tilde{\beta}_{u})$ if
$\tilde{\alpha}_{x}\leq\tilde{\beta}_{x},\qquad\tilde{\alpha}_{\theta}\leq\tilde{\beta}_{\theta},\qquad\tilde{\alpha}_{u}\leq\tilde{\beta}_{u},$
for some $\tilde{\beta}_{x},\tilde{\beta}_{\theta},\tilde{\beta}_{u}\leq 0$.
These exponents extend the analytic stability exponents of Section 4 to locally Lipschitz learning flows. Negative values correspond to asymptotic contractivity of the generalized linearised dynamics, while nonpositive values correspond to uniform boundedness of generalized sensitivities.
@@ 7.3 Variational energies and dissipation
We now extend the energy side of the theory. Since non smooth energies may not admit classical gradients, we work with proper lower semicontinuous functionals and first order difference inequalities. This follows nonsmooth Lyapunov theory for differential inclusions, proximal algorithms, and stochastic approximation [Clarke1990Optimization, Benaim1999, BolteSabachTeboulle2014, DavisDrusvyatskiy2019].
@@
$\mathcal{E}\colon\mathcal{X}_{0}\times\dots\times\mathcal{X}_{L}\times\Theta\to\mathbb{R}_{+}\cup\{+\infty\}$
that is finite on all admissible learning states and satisfies the coercivity bounds
$c_{1}\bigl(\|X_{L}\|^{2}+\|\theta\|^{2}\bigr)\leq\mathcal{E}(X_{0},\dots,X_{L},\theta)\leq c_{2}\bigl(\|X_{L}\|^{2}+\|\theta\|^{2}\bigr),$
@@
There exist a variational learning energy $\mathcal{E}$, a constant $\gamma>0$, and a perturbation remainder $R$ that is linear in $(\delta x,\delta\theta^{0},\delta u)$ at first order, such that for every training step $t$ and every admissible trajectory,
@@
Assumptions 5 and 6 are the non smooth analogues of Definition 3.3 and Assumption 2. Lower semicontinuity and coercivity ensure that $\mathcal{E}$ controls the norms of $X_{L}$ and $\theta$ even in the absence of differentiability, while the dissipation inequality encodes strict energy decay up to first order perturbations.
@@ 7.4 Fundamental Variational Stability Theorem
We now state the main result of the variational extension. It shows that the equivalence between bounded stability signatures and the existence of a dissipative energy persists in the non smooth regime.
@@
Let the flow maps $\Phi_{k,t}$ satisfy Assumption 4. Up to equivalence of norms on $\mathcal{X}_{0}\times\dots\times\mathcal{X}_{L}\times\Theta$, the following statements are equivalent.
@@
    Generalized differential stability. There exist constants $C_{x},C_{\theta},C_{u}\geq 1$ such that for all depths $L$, horizons $T$, and indices $0\leq k\leq L$, $0\leq t\leq T$,
@@
    Variational energy dissipative stability. There exists a variational learning energy $\mathcal{E}$ satisfying Assumption 5 and a constant $\gamma>0$ such that the dissipation inequality in Assumption 6 holds for all learning trajectories.
Moreover, whenever these equivalent conditions hold, the generalized stability indices $(\tilde{\beta}_{x},\tilde{\beta}_{\theta},\tilde{\beta}_{u})$ in Definition 7.3 can be chosen explicitly as functions of $(L^{\mathrm{C}}_{\mathrm{fwd}},L^{\mathrm{C}}_{\mathrm{par}},L^{\mathrm{C}}_{\mathrm{upd}})$ and $\gamma$.
Proof sketch.
The implication from (1) to (2) adapts the proof of Theorem 4.1. Uniform bounds on the generalized signatures imply incremental stability of the associated difference inclusion in a suitable product norm. Converse Lyapunov results for discrete time differential inclusions yield a regular Lyapunov function that decreases strictly along trajectories [Clarke1990Optimization, Clarke1998Nonsmooth]. By composing this function with the squared Euclidean norm of $(X_{L},\theta)$ one obtains a functional $\mathcal{E}$ satisfying Assumptions 5 and 6.
For the converse, coercivity of $\mathcal{E}$ controls $\|(X_{L},\theta)\|$ in terms of the energy. The dissipation inequality precludes unbounded growth of perturbations, since such growth would force the energy to increase along some subsequence. Selecting elements of $\partial_{C}\Phi_{k,t}(S^{t}_{k})$ and applying generalized chain rules [Clarke1990Optimization] yields uniform bounds on the linearised difference inclusion. These bounds imply uniform control of the generalized signatures and hence nonpositivity of the exponents. ∎
The theorem shows that analytic stability is preserved when passing from smooth to non smooth learning systems. The generalized exponents $(\tilde{\alpha}_{x},\tilde{\alpha}_{\theta},\tilde{\alpha}_{u})$ retain the same interpretation as in the smooth case: negative values correspond to contractivity of the learning flow, while nonpositive values correspond to global boundedness of sensitivities with respect to inputs, parameters, and update perturbations.
@@ 7.5 Examples: piecewise linear networks and proximal flows
We briefly illustrate how the variational framework applies to two representative non-smooth systems. In each case we indicate (i) uniform bounds on the generalized stability signatures and (ii) existence of a variational energy satisfying a discrete dissipation inequality. Detailed derivations follow the same pattern as in Sections 5, 6, and 7 and are therefore omitted.
@@
These examples show that architectural and optimisation-induced non-smoothness are handled within the same stability grammar by replacing classical Jacobians with Clarke derivatives and smooth energies with variational energies, while the generalized exponents and dissipation law retain the same form as in the smooth case.
@@ 8.1 Stability as a joint property of architecture and optimisation
In Sections 23, learning is modelled as a discrete-time dynamical system on a joint state space containing all layer representations and parameters. The Learning Stability Profile records operator norms of directional sensitivities derived from the Jacobian or generalized Jacobian of the coupled flow. The analytic and generalized exponents $(\alpha_{x},\alpha_{\theta},\alpha_{u})$ and $(\tilde{\alpha}_{x},\tilde{\alpha}_{\theta},\tilde{\alpha}_{u})$ summarise the asymptotic growth or decay of these sensitivities along depth and training time.
The Fundamental Analytic Stability Theorem 4.1 and the Fundamental Variational Stability Theorem 7.4 show that, in both smooth and non-smooth regimes, nonpositive exponents are equivalent to the existence of an energy functional satisfying a discrete dissipation inequality. The choice between smooth and variational formulations is dictated by regularity: $C^{1}$ systems admit a Jacobian-based analysis, while locally Lipschitz systems require Clarke generalized derivatives and lower semicontinuous energies. On their overlap, the two formulations coincide, since the Clarke generalized Jacobian reduces to the classical Jacobian on a dense set and the associated exponents agree up to norm equivalence [Clarke1990Optimization, RockafellarWets1998].
This joint perspective links local analytic quantities, such as spectral norms, Lipschitz constants, and step sizes, to global dynamical behaviour through stability exponents and energies. Architecture and optimisation are treated symmetrically as components of a single coupled flow. The feedforward, residual, stochastic gradient, and variational stability laws derived in Sections 5, 6, and 7 are specialisations of this unified criterion.
@@ 8.2 Connection with classical stability notions
The analytic–variational framework unifies stability concepts across several communities. In numerical analysis and control, Lyapunov theory relates contractivity and negative exponents to strict Lyapunov functions [Khalil2002Nonlinear], while CFL-type conditions ensure stability of explicit schemes for dissipative systems [Strikwerda2004]. In optimisation, firm nonexpansiveness and averagedness yield contractive behaviour for proximal and splitting methods [BauschkeCombettes2011, BolteSabachTeboulle2014, DavisDrusvyatskiy2019]. In statistical learning, algorithmic stability bounds generalisation error via insensitivity to perturbations [Bousquet2002, Hardt2016Train, BottouCurtisNocedal2018].
@@
The feedforward and residual stability laws recover discrete counterparts of stability results for neural ordinary differential equations [HaberRuthotto2017, RuthottoHaber2020]. The CFL-type step size conditions correspond to admissible time steps for explicit discretisations of dissipative flows, while the stochastic gradient stability law recovers classical stochastic approximation conditions [KushnerYin2003, Benaim1999] in exponent form.
@@ 8.3 Design rules and trade offs
For stochastic gradient methods, the temporal stability law expresses the parametric exponent in terms of learning rate, curvature, and gradient noise. The effective contraction factor $q(\eta)$ in Theorem 6.2 refines deterministic stability criteria by incorporating stochastic effects, explaining the sensitivity of training to learning-rate schedules in high-variance regimes [BottouCurtisNocedal2018].
The framework also clarifies accuracy–stability trade offs. Impossibility results for certain expressive architectures [ColbrookAntilTownsend2021] correspond to regimes where the forward exponent must become positive as accuracy is pushed beyond a threshold. The analytic exponents thus provide a principled way to express stability constraints in terms of geometry and spectral scale rather than empirical performance alone.
@@ 8.4 Robustness, generalisation, and explanations
The analytic and generalized exponents connect stability to robustness and generalisation. Algorithmic stability theory links small generalisation gaps to insensitivity of learning procedures [Bousquet2002, Hardt2016Train]. In the present framework, bounded parametric and temporal exponents imply that perturbations in training data or stochasticity induce uniformly bounded changes in parameters and representations, as quantified by the Learning Stability Profile.
The same mechanism governs the stability of explanation methods. Explanation functionals based on gradients or saliency inherit their sensitivity from network Jacobians and from the stability of the training procedure [Fel2022]. When forward or parametric exponents are close to zero or positive, small perturbations may be amplified, leading to unstable explanations even when prediction error is low. Negative exponents instead imply quantitative robustness of both predictions and a broad class of gradient-based explanations. While interpretability also involves fidelity and domain alignment [Abusitta2024], the analytic–variational framework isolates stability as a structural property governed by architecture and optimisation.
@@ 8.5 Scope and limitations
They do not cover all regimes encountered in practice. Strongly adaptive optimisers, non-stationary data, heavy-tailed noise, or unbounded parameter growth may violate global bounds or admit only local or state-dependent energies [BottouCurtisNocedal2018]. The theory is therefore structural rather than fully quantitative: the derived inequalities relate exponents, regularity constants, and dissipation rates, but are not claimed to be sharp in high-dimensional settings.
@@ 9 Conclusion
This paper develops an analytic and variational framework for stability in deep learning systems. The central object is the Learning Stability Profile, which records the differential response of coupled representation, parameter, and update dynamics. The Fundamental Analytic Stability Theorem and its variational extension show that uniform bounds on these sensitivities are equivalent to the existence of Lyapunov-type energies dissipating along the joint learning flow, without requiring convexity, stationarity, or architectural homogeneity beyond stated regularity assumptions.
Within this framework, explicit stability laws are derived for standard model classes. Feedforward networks admit spectral conditions for forward contractivity, residual architectures satisfy discrete CFL-type constraints linking step size and Jacobian bounds, and stochastic gradient methods exhibit temporal stability governed by curvature, noise, and learning rate. In each case, stability is encoded by a small set of exponents and dissipation rates directly traceable to architectural and optimisation choices.
The framework extends naturally to non-smooth systems through Clarke generalized derivatives and variational energy methods. Piecewise linear networks, proximal and projected updates, and stochastic subgradient flows admit generalized stability signatures and Lyapunov energies satisfying the same dissipation law as in the smooth case. The resulting exponents capture contractivity even in the presence of kinks, projections, and non-smooth updates.
Several directions remain open. Continuous-time limits would connect the discrete exponents to Lyapunov exponents of neural ordinary differential equations [HaberRuthotto2017, RuthottoHaber2020]. Extensions to curved parameter or representation manifolds would introduce curvature-dependent stability indices and link the analysis to information geometry. Further work is needed to relate analytic and generalized exponents to empirical training behaviour in large-scale models and to develop practical diagnostics for estimating stability indices from data.
In conclusion, deep learning systems can be analysed as discrete dynamical systems governed by verifiable energy laws. The Learning Stability Profile provides a structural language that turns heuristic stabilisation rules into precise conditions on the joint learning flow, offering a foundation for principled architecture and optimisation design with explicit analytic stability guarantees.
Appendix A Detailed Proofs
This appendix presents proofs omitted from the main text.
A.1 Proofs for Section 3
Proof of Proposition 3.4.
Let $|\cdot|_{1}$ and $|\cdot|_{2}$ be two norms on the finite-dimensional space $\mathcal{M}$. By norm equivalence, there exist constants $m,M>0$ such that
$m|v|_{1}\leq|v|_{2}\leq M|v|_{1}\qquad\text{for all }v\in\mathcal{M}.$ (A.1)
For any linear operator $A:\mathcal{M}\to\mathcal{M}$, the induced operator norms satisfy the standard equivalence bounds
$|A|_{2,\mathrm{op}}\leq\tfrac{M}{m}|A|_{1,\mathrm{op}},\qquad|A|_{1,\mathrm{op}}\leq\tfrac{M}{m}|A|_{2,\mathrm{op}},$ (A.2)
see, e.g., [HornJohnson2012, Section 5.4].
Applying these inequalities to the directional sensitivity operators $\mathsf{S}^{x}_{k,t}$ yields constants $c=m/M$ and $C=M/m$ such that
$c\,\sigma^{x,1}_{k,t}\leq\sigma^{x,2}_{k,t}\leq C\,\sigma^{x,1}_{k,t}\quad\text{for all }k,t.$
Taking finite-horizon suprema and dividing by $L+T$ gives additive constants $\tfrac{1}{L+T}\log c$ and $\tfrac{1}{L+T}\log C$, which vanish as $L+T\to\infty$. Hence the limsup defining $\alpha_{x}$ is invariant under the choice of norm. The same argument applies verbatim to $\mathsf{S}^{\theta}_{k,t}$ and $\mathsf{S}^{u}_{k,t}$, proving the claim. ∎
Proof of Proposition 3.5.
We treat the forward component; the parameter and temporal components follow identically.
(i) Uniform boundedness implies $\alpha_{x}\leq 0$. If $\sigma^{x}_{k,t}\leq C_{x}$ for all $k,t$, then the finite-horizon supremum $M_{L,T}\leq C_{x}$, and hence
$\alpha_{x}=\limsup_{L,T\to\infty}\frac{1}{L+T}\log M_{L,T}\leq\lim_{L,T\to\infty}\frac{\log C_{x}}{L+T}=0.$
(ii) $\alpha_{x}<0$ implies exponential decay of finite-horizon suprema. If $\alpha_{x}<0$, then by definition of the limsup there exists $\varepsilon>0$ and $N\in\mathbb{N}$ such that
$\frac{1}{L+T}\log M_{L,T}\leq-\varepsilon\quad\text{whenever }L+T\geq N.$ (A.3)
Rearranging yields
$M_{L,T}\leq e^{-\varepsilon(L+T)}\quad\text{for all }L+T\geq N,$
which proves the claimed exponential decay. The same reasoning applies to $\alpha_{\theta}$ and $\alpha_{u}$. ∎

### v2

@@ 6.1 Locally Lipschitz learning flows
is locally Lipschitz. There exist constants
@@
the corresponding blocks satisfy
@@
Local Lipschitz continuity guarantees that the Clarke generalized Jacobian is a nonempty compact convex set and varies upper semicontinuously Clarke [1990]. This assumption covers common non-smooth components used in learning systems, including piecewise affine activations, proximal operators, and convex projections Rockafellar and Wets [1998]; Bauschke and Combettes [2017].
Under Assumption 2, the learning dynamics are described by the difference inclusion
@@
with existence of trajectories ensured by standard results for discrete-time locally Lipschitz systems Clarke [1990].
6.2 Generalized Learning Stability Profile
We extend the Learning Stability Profile by measuring sensitivities using Clarke generalized Jacobians.
@@ 6.2 Generalized Learning Stability Profile
Under Assumption 2, the generalized stability signature at $(k,t)$ and state $S^{t}_{k}$ is
@@
When $\Phi_{k,t}$ is differentiable at $S^{t}_{k}$, the generalized profile coincides with the smooth Learning Stability Profile. In general, it captures worst-case local amplification over all active linearizations of the non-smooth learning flow.
6.3 Generalized analytic exponents
@@
$\tilde{\alpha}_{x}$ $:=\limsup_{L,T\to\infty}\frac{1}{L+T}\log\sup_{0\leq k\leq L,\;0\leq t\leq T}\sigma^{x,C}_{k,t},$
$\tilde{\alpha}_{\theta}$ $:=\limsup_{L,T\to\infty}\frac{1}{L+T}\log\sup_{0\leq k\leq L,\;0\leq t\leq T}\sigma^{\theta,C}_{k,t},$
$\tilde{\alpha}_{u}$ $:=\limsup_{L,T\to\infty}\frac{1}{L+T}\log\sup_{0\leq k\leq L,\;0\leq t\leq T}\sigma^{u,C}_{k,t}.$
@@
The learning system is generalized analytically stable if
Negative generalized exponents correspond to contractive generalized linearized dynamics, while nonpositive values correspond to uniform boundedness of sensitivities.
@@ 6.4 Variational energies and dissipation
$\mathcal{E}:\mathcal{X}_{0}\times\dots\times\mathcal{X}_{L}\times\Theta\to\mathbb{R}_{+}\cup\{+\infty\},$
finite on all admissible states and satisfying coercivity bounds
$c_{1}\bigl(\|X_{L}\|^{2}+\|\theta\|^{2}\bigr)\leq\mathcal{E}\leq c_{2}\bigl(\|X_{L}\|^{2}+\|\theta\|^{2}\bigr),$
@@
There exist a variational learning energy $\mathcal{E}$, a constant $\gamma>0$, and a perturbation remainder $R$, linear at first order in $(\delta x,\delta\theta^{0},\delta u)$, such that for every training step $t$,
@@ 6.5 Fundamental Variational Stability Theorem
Let the flow maps $\Phi_{k,t}$ satisfy Assumption 2. Up to equivalence of norms on $\mathcal{X}_{0}\times\dots\times\mathcal{X}_{L}\times\Theta$, the following are equivalent.
@@
    There exist constants $C_{x},C_{\theta},C_{u}\geq 1$ such that for all $k,t$,
@@
    There exists a variational learning energy $\mathcal{E}$ satisfying Assumptions 3 and 4.
Sketch of proof.
Uniform bounds on generalized stability signatures imply incremental stability of the associated difference inclusion under a suitable product norm. Converse Lyapunov results for nonsmooth discrete-time systems yield a coercive variational energy satisfying the dissipation inequality Clarke [1990]; Clarke et al. [1998].
Conversely, coercivity and dissipation preclude unbounded growth of generalized perturbations. Selecting elements of the Clarke generalized Jacobian and applying generalized chain rules yields uniform bounds on generalized sensitivities, implying nonpositivity of the generalized analytic exponents. ∎
The theorem shows that analytic stability is preserved when passing from smooth to non-smooth learning systems. Generalized analytic exponents retain the same interpretation as in the smooth case, quantifying contractivity or boundedness of learning dynamics with respect to inputs, parameters, and update mechanisms.
@@ 6.6 Examples: piecewise linear networks and proximal flows
We illustrate how the variational framework applies to two representative non-smooth systems by verifying uniform bounds on generalized stability signatures and the existence of a dissipative variational energy. Detailed derivations follow the same pattern as in Sections 4, 5, and 6 and are therefore omitted.
@@
These examples illustrate that architectural and optimisation-induced non-smoothness fit naturally within the same stability grammar.
@@ 7.1 Stability as a joint property of architecture and optimisation
In previous sections, learning is modelled as a discrete-time dynamical system on a joint state space containing all layer representations and parameters. The Learning Stability Profile records operator norms of directional sensitivities derived from the Jacobian or generalized Jacobian of the coupled flow. The analytic and generalized exponents $(\alpha_{x},\alpha_{\theta},\alpha_{u})$ and $(\tilde{\alpha}_{x},\tilde{\alpha}_{\theta},\tilde{\alpha}_{u})$ summarise the asymptotic growth or decay of these sensitivities along depth and training time.
The Fundamental Analytic Stability Theorem 3.1 and the Fundamental Variational Stability Theorem 6.4 show that, in both smooth and non-smooth regimes, nonpositive exponents are equivalent to the existence of an energy functional satisfying a discrete dissipation inequality. The choice between smooth and variational formulations is dictated by regularity: $C^{1}$ systems admit a Jacobian-based analysis, while locally Lipschitz systems require Clarke generalized derivatives and lower semicontinuous energies. On their overlap, the two formulations coincide, since the Clarke generalized Jacobian reduces to the classical Jacobian on a dense set and the associated exponents agree up to norm equivalence Clarke [1990]; Rockafellar and Wets [1998].
This joint perspective links local analytic quantities, such as spectral norms, Lipschitz constants, and step sizes, to global dynamical behaviour through stability exponents and energies. Architecture and optimisation are treated symmetrically as components of a single coupled flow. The feedforward, residual, stochastic gradient, and variational stability laws derived in Sections 4, 5, and 6 are specialisations of this unified criterion.
@@ 7.2 Connection with classical stability notions
The analytic–variational framework unifies stability concepts across several communities. In numerical analysis and control, Lyapunov theory relates contractivity and negative exponents to strict Lyapunov functions Khalil [2002], while CFL-type conditions ensure stability of explicit schemes for dissipative systems Strikwerda [2004]. In optimisation, firm nonexpansiveness and averagedness yield contractive behaviour for proximal and splitting methods Bauschke and Combettes [2017]; Bolte et al. [2014]; Davis and Drusvyatskiy [2019]. In statistical learning, algorithmic stability bounds generalisation error via insensitivity to perturbations Bousquet and Elisseeff [2002]; Hardt et al. [2016]; Bottou et al. [2018].
@@
The feedforward and residual stability laws recover discrete counterparts of stability results for neural ordinary differential equations Haber and Ruthotto [2017]; Ruthotto and Haber [2020]. The CFL-type step size conditions correspond to admissible time steps for explicit discretisations of dissipative flows, while the stochastic gradient stability law recovers classical stochastic approximation conditions Kushner and Yin [2003]; Benaïm [1999] in exponent form.
@@ 7.3 Design rules and trade offs
For stochastic gradient methods, the temporal stability law expresses the parametric exponent in terms of learning rate, curvature, and gradient noise. The effective contraction factor $q(\eta)$ in Theorem 5.2 refines deterministic stability criteria by incorporating stochastic effects, explaining the sensitivity of training to learning-rate schedules in high-variance regimes Bottou et al. [2018].
The framework also clarifies accuracy–stability trade offs. Impossibility results for certain expressive architectures Colbrook et al. [2021] correspond to regimes where the forward exponent must become positive as accuracy is pushed beyond a threshold. The analytic exponents thus provide a principled way to express stability constraints in terms of geometry and spectral scale rather than empirical performance alone.
@@ 7.4 Robustness, generalisation, and explanations
The analytic and generalized exponents connect stability to robustness and generalisation. Algorithmic stability theory links small generalisation gaps to insensitivity of learning procedures Bousquet and Elisseeff [2002]; Hardt et al. [2016]. In the present framework, bounded parametric and temporal exponents imply that perturbations in training data or stochasticity induce uniformly bounded changes in parameters and representations, as quantified by the Learning Stability Profile.
The same mechanism governs the stability of explanation methods. Explanation functionals based on gradients or saliency inherit their sensitivity from network Jacobians and from the stability of the training procedure Fel et al. [2022]. When forward or parametric exponents are close to zero or positive, small perturbations may be amplified, leading to unstable explanations even when prediction error is low. Negative exponents instead imply quantitative robustness of both predictions and a broad class of gradient-based explanations. While interpretability also involves fidelity and domain alignment Abusitta et al. [2024], the analytic–variational framework isolates stability as a structural property governed by architecture and optimisation.
@@ 7.5 Scope and limitations
They do not cover all regimes encountered in practice. Strongly adaptive optimisers, non-stationary data, heavy-tailed noise, or unbounded parameter growth may violate global bounds or admit only local or state-dependent energies Bottou et al. [2018]. The theory is therefore structural rather than fully quantitative: the derived inequalities relate exponents, regularity constants, and dissipation rates, but are not claimed to be sharp in high-dimensional settings.
@@ 8 Conclusion
We introduced an analytic and variational framework for stability in deep learning systems based on the Learning Stability Profile, which quantifies how perturbations propagate through coupled representation and parameter dynamics. The Fundamental Analytic Stability Theorem and its variational extension show that bounded sensitivities are equivalent to the existence of dissipative Lyapunov-type energies, without requiring convexity or algorithm-specific structure.
Within this framework, classical stability conditions for feedforward, residual, and stochastic gradient methods arise as special cases, while the theory extends naturally to non-smooth systems through Clarke generalized derivatives and variational energies. Stability is therefore expressed through a small set of exponents directly linked to architectural and optimization choices.
Future directions include continuous-time limits, geometric formulations of learning dynamics, and practical estimation of stability indices in large-scale models. Overall, the framework provides a structural basis for principled architecture and optimization design grounded in analytic stability guarantees.
Appendix A Technical Proofs
This appendix contains proofs of results stated in the main text. Standard facts repeatedly used include equivalence of norms in finite dimensions, operator norm bounds, and classical Lyapunov arguments; see [Horn and Johnson, 2013, Section 5.4] and [Khalil, 2002, Chapter 4].
Only arguments specific to the present framework are detailed. Classical steps are cited rather than reproved.
A.1 Proof of Theorem 3.1
Proof.
Let
be equipped with an arbitrary norm. All norms on $\mathcal{M}$ are equivalent.
Denote by $\Psi_{t}$ the one-step learning map $S^{t}_{\cdot}\mapsto S^{t+1}_{\cdot}$ obtained by composing depthwise representation updates with parameter updates at time $t$.
[…]
Uniform bounds on the signatures $\sigma^{x}_{k,t}$, $\sigma^{\theta}_{k,t}$, and $\sigma^{u}_{k,t}$ imply uniform bounds on derivatives of $\Psi_{t}$ after passing to an equivalent product norm. Hence each $\Psi_{t}$ is globally Lipschitz.
Nonpositive analytic exponents mean perturbations do not grow uniformly over depth and time, so the coupled dynamics are incrementally stable. Converse Lyapunov theorems for incrementally stable discrete systems then yield a $C^{1}$ incremental Lyapunov function Angeli [2002]; Khalil [2002] satisfying quadratic bounds and strict decrease along trajectories.
Fixing a reference trajectory converts this incremental Lyapunov function into a coercive learning energy $\mathcal{E}$ whose one-step variation satisfies the dissipation inequality of item (2). First-order perturbation terms produce the remainder $R$.
[…]
Assume a coercive differentiable energy satisfies the dissipation inequality. Setting perturbations to zero yields strict decay along trajectories. Coercivity bounds state norms of relevant coordinates are controlled by $\mathcal{E}$, so growth of perturbations would contradict energy decay.
Linearizing the dynamics yields a linear time-varying system whose transition operators must therefore remain uniformly bounded [Khalil, 2002, Chapter 4]. Projection onto coordinate blocks gives uniform bounds on the sensitivity operators, hence $\alpha_{x},\alpha_{\theta},\alpha_{u}\leq 0$.
Quantitative index bounds follow from quadratic comparison constants and norm equivalence. ∎